OpenStudy (anonymous):

Evaluate the expression 5log(x-3) do not use a calculator

OpenStudy (lgbasallote):

no given value for x?

OpenStudy (anonymous):

NOPE

OpenStudy (anonymous):

THis is stupid.

hero (hero):

Good luck with that

OpenStudy (anonymous):

would I just graph it?

hero (hero):

Says "evaluate" not graph. But I don't see how it is possible to evaluate without knowing what x equals

OpenStudy (anonymous):

\[5\log(x-3) = \log(x-3)^5\] this is wat u can do without the value of x

OpenStudy (anonymous):

A Mathematica plot of \[\left\{5 \log (x-3),\log ^5(x-3)\right\} \]in blue and red respectively, is attached.